# MIBench: Evaluating Multimodal Large Language Models over Multiple Images

### v1

@@ Abstract
Built on the power of LLMs, numerous multimodal large language models (MLLMs) have recently achieved remarkable performance on various vision-language tasks across multiple benchmarks. However, most existing MLLMs and benchmarks primarily focus on single-image input scenarios, leaving the performance of MLLMs when handling realistic multiple images underexplored. Although a few benchmarks consider multiple images, their evaluation dimensions and samples are very limited. Therefore, in this paper, we propose a new benchmark MIBench, to comprehensively evaluate fine-grained abilities of MLLMs in multi-image scenarios. Specifically, MIBench categorizes the multi-image abilities into three scenarios: multi-image instruction (MII), multimodal knowledge-seeking (MKS) and multimodal in-context learning (MIC), and constructs 13 tasks with a total of 13K annotated samples. During data construction, for MII and MKS, we extract correct options from manual annotations and create challenging distractors to obtain multiple-choice questions. For MIC, to enable an in-depth evaluation, we set four sub-tasks and transform the original datasets into in-context learning formats. We evaluate several open-source MLLMs and close-source MLLMs on the proposed MIBench. The results reveal that although current models excel in single-image tasks, they exhibit significant shortcomings when faced with multi-image inputs, such as confused fine-grained perception, limited multi-image reasoning, and unstable in-context learning. The annotated data in MIBench is available at https://huggingface.co/datasets/StarBottle/MIBench.
@@
Haowei Liu1,2, Xi Zhang3, Haiyang Xu3†, Yaya Shi4, Chaoya Jiang5, Ming Yan3, Ji Zhang3, Fei Huang3, Chunfeng Yuan1,2†, Bing Li1,2, Weiming Hu1,6 1MAIS, Institute of Automation, Chinese Academy of Sciences, China 2School of Artificial Intelligence, University of Chinese Academy of Sciences, China 3Alibaba Group 4University of Science and Technology of China 5Peking University 6School of Information Science and Technology, ShanghaiTech University, China liuhaowei2019@ia.ac.cn, cfyuan@nlpr.ia.ac.cn {shuofeng.xhy, ym119608}@alibaba-inc.com
@@ 1 Introduction
Recently, leveraging the powerful comprehension and reasoning abilities of LLMs, many MLLMs (such as LLaVA-1.5(Liu et al., 2024) and mPLUG-Owl2(Ye et al., 2024)) have demonstrated outstanding performance across various vision-language tasks (e.g. image captioning, VQA and visual grounding). Concurrently, numerous benchmarks (e.g. MME(Fu et al., 2023), MMBench(Liu et al., 2023) and SEED-Bench(Li et al., 2024)) have been proposed to evaluate the abilities of MLLMs in terms of different perspectives such as recognition, localization and reasoning.
We evaluate several existing MLLMs on the proposed MIBench, including both closed-source (e.g. GPT-4V) and open-source models (e.g. LLaVA-1.5, Qwen-VL and IDEFICS2). The evaluation results reveal that current MLLMs especially open-source models have major flaws in multi-image scenarios. We’ll release our benchmark to spur progress in improving the multi-image abilities of MLLMs.
Our contributions are summarized as follows:
  • We propose the first large-scale and comprehensive benchmark MIBench for evaluating the multi-image abilities of MLLMs, covering three scenarios and 13 tasks.
  • Current MLLMs perform poorly in the multimodal knowledge-seeking scenario, due to limited abilities of multi-image perception and comparison. And there remains considerable room for improvement in multimodal in-context learning abilities.
@@ 2.1 Multimodal Large Language Models
Recent research (Zhu et al., 2023; Liu et al., 2024; Dai et al., 2024; Ye et al., 2023) has expanded LLMs (e.g. LLaMA Touvron et al., 2023) into multimodal scenarios, enabling them to process both visual and textual information. Some studies (Jiang et al., 2024; Huang et al., 2023; Laurençon et al., 2024) have further explored augmenting MLLMs with multi-image understanding ability. However, there lacks a comprehensive benchmark for multi-image understanding abilities, which limits the full exploration of these models’ potential and hinders the development of this field.
@@ 2.2 MLLM Benchmarks
The rapid development of MLLMs has led to the emergence of a series of benchmarks, like LVLM-eHub (Xu et al., 2023), SEED-bench (Li et al., 2023a), MMBench (Liu et al., 2023), and MM-Vet (Yu et al., 2023). However, these benchmarks primarily focus on single-image evaluation, and often overlook multi-image perception and reasoning abilities, which hold greater practical value.
Recent studies have developed benchmarks for assessing multi-image reasoning capabilities. SparklesEval aims to establish a benchmark for multi-turn dialogues and multi-image scenarios. However, it exclusively focuses on dialogue scenario, and relies entirely on GPT-4 for evaluation. Besides, its scale is small. Other datasets such as Mantis-Eval (Jiang et al., 2024) and SEED-Bench2 (Li et al., 2024) also cover a limited number of multi-image tasks and their scale is constrained due to reliance on manual annotations.
In this paper, we propose a large-scale benchmark with 3 scenarios and 13 tasks, comprising 13K high-quality annotations, to comprehensively evaluate the multi-image capabilities of MLLMs.
@@ 3 MIBench
We categorize the multi-image inputs into three scenarios: Multi-Image Instruction (MII), Multimodal Knowledge-Seeking (MKS) and Multimodal In-Context Learning (MIC). As Figure 2 shows, MII refers to cases where instructions involve perception, comparison and reasoning across multiple images. For instance, "Do the two images show the same number of cats?" In contrast, MKS provides models with interleaved image-text data as external knowledge, while the question itself is about an image or independent of any image. MIC is another popular scenario, in which MLLMs respond to visual questions while being provided with a series of multimodal demonstrations (i.e., demos).
@@ 3.1.1 Multi-Image Instruction
Logical Reasoning (LR) task requires the model to perform logical reasoning and analyze the causal relationships between objects or events shown in the images. The video QA dataset NExT-QA (Xiao et al., 2021) is used for data construction.
@@ 3.1.2 Multimodal Knowledge-Seeking
Text-Rich Images (TRI) VQA task evaluates the model’s ability to understand text-rich images and extract information relevant to the given question, which is very common in real-world scenarios (e.g. reading slides or documents). We adopt the SlideVQA (Tanaka et al., 2023) dataset for data construction.
Vision-linked Textual Knowledge (VTK) task corresponds to a very practical scenario where the question is beyond the visual content of the query image, such as querying background knowledge. The provided outside knowledge encompasses images and corresponding text possibly retrieved from a knowledge base (e.g., Wikipedia). The model is required to link the query image to the relevant image, and extract useful information from the corresponding text. Figure 2(h) shows an example, whose data is from the InfoSeek dataset (Chen et al., 2023).
Text-linked Visual Knowledge (TVK) task refers to cases where the text-only question is about the visual attributes of a specific object. For instance, "Is the China National Stadium round or square?". When provided with outside knowledge in an interleaved image-text form, the model needs to link the question to the relevant text, and extract visual information from the corresponding image. This task is very common in real life such as browsing web pages. Figure 2(i) shows an example, whose data is from the WebQA dataset (Chang et al., 2022).
@@ 3.1.3 Multimodal In-Context Learning
The in-context learning ability enables LLMs to gain performance boost when provided with a series of demos. Flamingo (Alayrac et al., 2022), OpenFlamingo (Awadalla et al., 2023) and IDEFICS (Laurençon et al., 2024) have explored multimodal in-context learning (MIC). For the evaluation of the MIC ability, existing methods solely assess the model’s performance via a holistic metric, such as accuracy on the VQAv2 (Goyal et al., 2017b) dataset. To evaluate the model’s MIC ability in a fine-grained manner, we categorize the MIC scenario into the following four distinct tasks:
@@
Open-ended VQA task has an open range of possible answers which cannot be fully covered by the demos. The task evaluates the model’s ability to learn task patterns through demos. We conduct a balanced sampling of different knowledge types of the OK-VQA dataset (Marino et al., 2019) for this task.
Hallucination phenomenon is a significant challenge faced by MLLMs. In this task, we construct the hallucination dataset POPE (Li et al., 2023b) into in-context learning format, and study the impact of the model’s MIC ability on the hallucination phenomenon.
Demo-based Task Learning is a core aspect of in-context learning, which enables the model to rapidly adapt to new tasks given a few demos. To investigate existing MLLMs’ demo-based task learning ability, we select several visual tasks from the VQAv2 dataset and remove the task instructions. Instead, we present the task demos in the form like "rabbit: 3". Figure 2(m) shows an example.
@@ 3.2 Data Generation
In Section 3.1, we introduced the evaluation tasks and the corresponding data source of the proposed MIBench. However, the generation of test samples using the original annotations is nontrivial. We meticulously devise a data generation pipeline, including various strategies of question generation, distractor generation and external knowledge sampling for different tasks.
Question Generation. To enhance the diversity of questions, we devise corresponding prompts for each task, and employ GPT-4 to generate a variety of question forms. We then randomly sample from the question pool to construct the test samples. For instance, for the General Comparison (GC) task, the questions such as "Is the subsequent sentence an accurate portrayal of the two images?" and "Can the given sentence accurately illustrate what’s in these two images?" are utilized.
Distractor Generation. For different tasks, we adopt two methods of distractor generation. One way is to sample from the original annotations following certain strategies. For instance, for the Temporal Reasoning (TR) task, we utilize the Something-something V2 dataset for data construction. To prevent the model from taking shortcuts by identifying objects to choose the correct options, we sample different temporal relationships of the same object from the annotations as distractors. In this way, the constructed test samples can more accurately reflect the model’s understanding of temporal relationships. The second method is to generate distractors with the help of GPT-4. For instance, in the Text-Rich Images (TRI) task, we prompt GPT-4 to generate distractors according to the question and the correct answer.
External Knowledge Sampling. For the Multimodal Knowledge-Seeking (MKS) scenario, reasonably sampling interleaved image-text data as external knowledge is very important to the quality of test samples. For instance, in the Vision-linked Textual Knowledge (VTK) task, we select text and images from the original annotations which have the same question as the current query but with different answers as external knowledge. This approach avoids selecting text and images unrelated to the current query, and can thus generate more challenging distractors.
@@ 3.3 Quality Control
To mitigate data contamination, our construction of test data exclusively utilizes the validation or test sets from existing datasets. Furthermore, to ensure the high quality and reliability of the test data, we combine automated filtering and manual verification to remove samples with ambiguous or duplicate options.
Moreover, in the Multimodal Knowledge-Seeking scenario, our goal is to assess the model’s ability to utilize external knowledge to answer questions, rather than relying on its inherent knowledge. Therefore, to prevent the varying inherent knowledge of different models from impairing the evaluation, we conduct pre-tests using multiple MLLMs and eliminate those samples which the models can correctly answer without external knowledge.
@@ 3.4 Evaluation
For the multiple-choice questions, we employ the accuracy of the predicted options as the evaluation metric. Notably, early MLLMs such as mPLUG-Owl tend to produce longer responses rather than directly outputting the option. For these models, we use GPT-4 to determine which option matches the predicted content. In addition, similar to the observation of MMBench, we find that different MLLMs show preferences for specific options (i.e. position bias). Therefore, we set the correct option sequentially to ’A’, ’B’, ’C’ and ’D’. A model is considered to have correctly answered a sample only if it consistently provides the correct response across multiple tests. In this way, the impact of position bias on the evaluation results is mitigated.
@@ 4.1 Models
In this section, we evaluate MLLMs using the constructed MIBench dataset. We first evaluate MLLMs on the Multi-Image INstruction and Multimodal Knowledge-Seeking scenarios. These models can be categorized into three distinct groups: (1) close-source models, including GPT-4V and GPT-4o; (2) open-source single-image MLLMs, including mPLUG-Owl, LLaVA-1.5, Qwen-VL and mPLUG-Owl2; (3) open-source models natively supporting multi-image input, including Mantis and IDEFICS2. For the open-source models, we employ greedy decoding for prediction generation.
Then we evaluate MLLMs on the Multimodal In-Context Learning scenario. Currently, apart from OpenFlamingo, MMICL and IDEFICS, most MLLMs have not explored multimodal ICL abilities. As a result, they exhibit consistent negative ICL effect in testing, i.e. the increase of demos leads to performance decline. Therefore, we only present the results of models that possess multimodal ICL abilities.

### v2

@@ Abstract
Built on the power of LLMs, numerous multimodal large language models (MLLMs) have recently achieved remarkable performance on various vision-language tasks. However, most existing MLLMs and benchmarks primarily focus on single-image input scenarios, leaving the performance of MLLMs when handling realistic multiple images underexplored. Although a few benchmarks consider multiple images, their evaluation dimensions and samples are very limited. In this paper, we propose a new benchmark MIBench, to comprehensively evaluate fine-grained abilities of MLLMs in multi-image scenarios. Specifically, MIBench categorizes the multi-image abilities into three scenarios: multi-image instruction (MII), multimodal knowledge-seeking (MKS) and multimodal in-context learning (MIC), and constructs 13 tasks with a total of 13K annotated samples. During data construction, for MII and MKS, we extract correct options from manual annotations and create challenging distractors to obtain multiple-choice questions. For MIC, to enable an in-depth evaluation, we set four sub-tasks and transform the original datasets into in-context learning formats. We evaluate several open-source and closed-source MLLMs on the proposed MIBench. The results reveal that although current models excel in single-image tasks, they exhibit significant shortcomings when faced with multi-image inputs, such as limited fine-grained perception, multi-image reasoning and in-context learning abilities. The annotated data of MIBench is available at https://huggingface.co/datasets/StarBottle/MIBench.
@@
Haowei Liu1,2, Xi Zhang3, Haiyang Xu3†, Yaya Shi4, Chaoya Jiang5, Ming Yan3, Ji Zhang3, Fei Huang3, Chunfeng Yuan1,2†, Bing Li1,2, Weiming Hu1,2,6 1MAIS, Institute of Automation, Chinese Academy of Sciences, China 2School of Artificial Intelligence, University of Chinese Academy of Sciences, China 3Alibaba Group 4University of Science and Technology of China 5Peking University 6School of Information Science and Technology, ShanghaiTech University, China liuhaowei2019@ia.ac.cn, cfyuan@nlpr.ia.ac.cn {shuofeng.xhy, ym119608}@alibaba-inc.com
@@ 1 Introduction
Recently, leveraging the powerful comprehension and reasoning abilities of LLMs, many MLLMs such as LLaVA-1.5 (Liu et al., 2024) and mPLUG-Owl2 (Ye et al., 2024b) have demonstrated outstanding performance across various vision-language tasks (e.g. image captioning, VQA and visual grounding). Concurrently, numerous benchmarks like MME (Fu et al., 2023), MMBench (Liu et al., 2023) and SEED-Bench (Li et al., 2024)) have been proposed to evaluate the abilities of MLLMs in terms of different perspectives such as recognition, localization and reasoning.
We evaluate several existing MLLMs on the proposed MIBench, including both closed-source (e.g. GPT-4o) and open-source models (e.g. LLaVA-1.5, Idefics2 and mPLUG-Owl3). The evaluation results reveal that current MLLMs especially open-source models have major flaws in multi-image scenarios. The annotated data of our MIBench is publicly available to spur progress in improving the multi-image abilities of MLLMs.
Our contributions are summarized as follows:
  • We propose the first large-scale and comprehensive benchmark MIBench for evaluating the multi-image abilities of MLLMs, covering three scenarios and 13 tasks in total.
  • Current MLLMs perform poorly in the multimodal knowledge-seeking scenario. And there still exists considerable room for improvement in the multimodal in-context learning abilities.
@@ 2.1 Multimodal Large Language Models
Recent research (Zhu et al., 2023; Liu et al., 2024; Dai et al., 2024; Ye et al., 2023) has expanded LLMs (e.g. LLaMA Touvron et al., 2023) into multimodal scenarios, enabling them to process both visual and textual information. Some studies (Jiang et al., 2024; Huang et al., 2023; Laurençon et al., 2024; Ye et al., 2024a) have further explored augmenting MLLMs with multi-image understanding abilities. However, there lacks a comprehensive benchmark for evaluating the multi-image abilities of MLLMs, which limits the full exploration of these models’ potential and hinders the development of this field.
@@ 2.2 MLLM Benchmarks
The rapid development of MLLMs has led to the emergence of a series of benchmarks, such as LVLM-eHub (Xu et al., 2023), MMBench (Liu et al., 2023), MM-Vet (Yu et al., 2023) and SEED-Bench (Li et al., 2023a). However, these benchmarks primarily focus on single-image evaluation, and often overlook multi-image perception and reasoning abilities, which hold even greater practical value. Some recent studies develop benchmarks for assessing multi-image capabilities. Sparkles-Eval aims to establish a benchmark for multi-turn dialogues and multi-image scenarios. However, it exclusively focuses on the dialogue scenario, and relies entirely on GPT-4 for evaluation. Besides, it has a small data scale. Other datasets such as Mantis-Eval (Jiang et al., 2024) and SEED-Bench2 (Li et al., 2024) also cover a small number of multi-image tasks, with a limited scale due to reliance on manual annotation.
In this paper, we propose a large-scale benchmark covering three multi-image scenarios and 13 tasks, to comprehensively evaluate the multi-image capabilities of MLLMs.
@@ 3.1 Evaluation Taxonomy
We categorize multi-image inputs into three scenarios: Multi-Image Instruction (MII), Multimodal Knowledge-Seeking (MKS) and Multimodal In-Context Learning (MIC). As Figure 2 shows, MII refers to cases where instructions involve perception, comparison and reasoning across multiple images. For instance, “Do the two images show the same number of cats?” MKS examines the ability of MLLMs to acquire relevant information from external knowledge, which is provided in an interleaved image-text format. Compared to MII, the questions in the MKS scenario can be about a single image or even independent of any visual content. MIC is another popular scenario, in which MLLMs respond to visual questions while being provided with a series of multimodal demonstrations (i.e., demos).
@@ 3.1.1 Multi-Image Instruction
According to the semantic types of the instructions, we further categorize the Multi-Image Instruction scenario into the following five tasks.
@@
Logical Reasoning (LR) task requires the model to perform logical reasoning and analyze the causal relationships between objects or events shown in the input images. The video QA dataset NExT-QA (Xiao et al., 2021) is used for data construction.
@@ 3.1.2 Multimodal Knowledge-Seeking
Based on the forms of external knowledge, we categorize the Multimodal Knowledge-Seeking scenario into the following four tasks.
@@
Text-Rich Images (TRI) VQA task evaluates the model’s ability to understand text-rich images and extract information relevant to the question, which is very common in real-world scenarios (e.g. reading slides or documents). We adopt the SlideVQA (Tanaka et al., 2023) dataset for data construction.
Vision-linked Textual Knowledge (VTK) task corresponds to a very practical scenario where the question is beyond the visual content of the query image, such as querying background knowledge. The provided external knowledge encompasses images and corresponding text which are possibly retrieved from a knowledge base (e.g., Wikipedia). The model is required to link the query image to the relevant image, and extract useful information from the corresponding text. Figure 2(h) shows an example, whose data is from the InfoSeek dataset (Chen et al., 2023).
Text-linked Visual Knowledge (TVK) task refers to cases where the text-only question is about the visual attributes of a specific object. For instance, "Is the China National Stadium round or square?" When provided with external knowledge in an interleaved image-text form, the model needs to link the question to the relevant text, and extract visual information from the corresponding image. This task is very common in real life such as browsing web pages. Figure 2(i) shows an example, whose data is from the WebQA dataset (Chang et al., 2022).
@@ 3.1.3 Multimodal In-Context Learning
The in-context learning ability enables LLMs to gain performance boost when provided with a series of demos. Recent studies (Alayrac et al., 2022; Awadalla et al., 2023; Laurençon et al., 2024) have also explored multimodal in-context learning (MIC). For the evaluation of the MIC ability, existing methods solely assess the model’s performance via a holistic metric, such as accuracy on the VQAv2 (Goyal et al., 2017b) dataset. To evaluate the model’s MIC ability in a fine-grained manner, we categorize the MIC scenario into the following four distinct tasks.
@@
Open-ended VQA task has an open range of possible answers which cannot be fully covered by the provided demos. The task evaluates the model’s ability to learn task patterns through demos. We conduct a balanced sampling of different knowledge types from the OK-VQA dataset (Marino et al., 2019) for this task.
Hallucination phenomenon is a significant challenge faced by MLLMs. In this task, we convert the hallucination dataset POPE (Li et al., 2023b) into in-context learning format, and study the impact of the model’s MIC ability on the hallucination phenomenon.
Demo-based Task Learning is a core aspect of in-context learning, which enables the model to rapidly adapt to new tasks given a few demos. To investigate existing MLLMs’ demo-based task learning ability, we select several visual tasks from the VQAv2 dataset and remove the task instructions. Instead, we present the task demos in the form like “rabbit: 3”. Figure 2(m) shows an example.
@@ 3.2 Data Generation
In Section 3.1, we introduced the evaluation tasks and the corresponding data source of the proposed MIBench. However, the generation of test samples using the original datasets is nontrivial. We meticulously devise a data generation pipeline, including various strategies of question generation, distractor generation and external knowledge sampling for different tasks.
Question Generation. To enhance the diversity of questions, we devise corresponding prompts for the tasks, and employ GPT-4 to generate a variety of question forms. We then randomly sample from the question pool to construct the test samples. For instance, for the General Comparison (GC) task, the questions such as “Is the subsequent sentence an accurate portrayal of the two images?” and “Can the given sentence accurately illustrate what’s in these two images?” are utilized.
Distractor Generation. For different tasks, we adopt two methods of distractor generation. One way is to sample from the original annotations following certain strategies. For instance, for the Temporal Reasoning (TR) task, we utilize the Something-something V2 dataset for data construction. To prevent the model from taking shortcuts by identifying objects to choose the correct options, we sample different temporal relationships of the same object from the annotations as distractors. In this way, the constructed test samples can more accurately reflect the model’s understanding of temporal relationships. The second method is to generate distractors with the help of GPT-4. For instance, in the Text-Rich Images (TRI) VQA task, we prompt GPT-4 to generate distractors according to the question and the correct answer.
External Knowledge Sampling. For the Multimodal Knowledge-Seeking (MKS) scenario, reasonably sampling interleaved image-text data as external knowledge is very important to the quality of test samples. For instance, in the Vision-linked Textual Knowledge (VTK) task, we select text and images from the original annotations which have the same question as the current query but with different answers as external knowledge. This approach avoids selecting text and images unrelated to the current query, and can thus generate more challenging distractors. Additionally, some datasets require more complex information extraction. For instance, we use GPT-4 to extract question-related segments from the original wiki entries of the InfoSeek dataset, which can be as long as several thousand words.
@@ 3.3 Quality Control
To mitigate data contamination, our construction of test data exclusively utilizes the validation or test sets from existing datasets. Furthermore, we combine automated filtering and manual verification to ensure the quality and reliability of the test data.
Specifically, after the data generation process, we perform two automated filtering strategies on the obtained data. 1) We remove images from the input samples, and test multiple advanced MLLMs on them. Then we discard samples which can still be answered correctly without visual input. This avoids the overestimation of model performance due to the textual bias of the questions and options. 2) For the Multimodal Knowledge-Seeking scenario, we eliminate external knowledge from the samples and test them using multiple MLLMs. Then we remove samples which the models can answer correctly without external knowledge. This mitigates the impact of internal knowledge of the model, and provides a more accurate assessment of the model’s ability of utilizing external knowledge.
As stated in Section 3.2, for some tasks such as Visual Referring, we employ GPT-4 to generate distractors. To ensure the high quality of the generated samples, we apply manual verification after automated filtering. The process is conducted by three trained annotators who possess relevant professional backgrounds. Specifically, a sample is discarded if there are duplicate options or more than one correct option.
@@ 3.4 Evaluation
For the multiple-choice questions, we employ the accuracy of the predicted options as the evaluation metric. Notably, early MLLMs such as mPLUG-Owl tend to produce longer responses rather than directly outputting the option. For these models, we use GPT-4 to determine which option matches the predicted content. In addition, similar to the observation of MMBench, we find that different MLLMs show preferences for specific options (i.e. position bias). Therefore, we set the correct option sequentially to “A”, “B”, “C” and “D”. A model is considered to have correctly answered a sample only if it consistently provides the correct response across multiple tests. In this way, the impact of position bias on the evaluation results is mitigated.
@@ 4 Experiments
Figure 3: A qualitative case of the Subtle Difference task, where open-source MLLMs show inferior performance due to limited fine-grained perception ability.
@@ 4.1 Models
In this section, we evaluate MLLMs using the constructed MIBench dataset. We first evaluate MLLMs on the Multi-Image Instruction and Multimodal Knowledge-Seeking scenarios. These models can be categorized into three distinct groups: (1) closed-source models, including GPT-4V and GPT-4o; (2) open-source single-image MLLMs, including mPLUG-Owl (Ye et al., 2023), LLaVA-1.5 (Liu et al., 2024), Qwen-VL (Bai et al., 2023) and mPLUG-Owl2 (Ye et al., 2024b); (3) open-source models natively supporting multi-image input, including Mantis (Jiang et al., 2024), Idefics2 (Laurençon et al., 2024) and mPLUG-Owl3 (Ye et al., 2024a). For the open-source models, we employ greedy decoding for prediction generation.
Then we evaluate open-source MLLMs on the Multimodal In-Context Learning (MIC) scenario. However, as most of these models have neither been pre-trained on large-scale interleaved image-text data nor fine-tuned on ICL format data, they do not exhibit MIC capabilities. In the tests across the four MIC tasks, they consistently exhibit a negative ICL effect, i.e., their performance decreases as the number of demos increases. Therefore, we only present the evaluation results of models that possess multimodal ICL abilities, i.e. OpenFlamingo, MMICL, IDEFICS and IDEFICS-I.